OpenStudy (anonymous):

Player 3-Point Shots Field Goals Free Throws Williams 2 7 6 Porter 5 4 5 Three-point shots are worth 3 points each. Field goals are worth 2 points each. Free throws are worth 1 point each. How many points did each player score? Use matrices to solve the problem.

OpenStudy (anonymous):

@RoseDryer

OpenStudy (anonymous):

\[\left[\begin{matrix}2 & 7 & 6 \\ 5 & 4\ & 5\end{matrix}\right]\] First matix.

OpenStudy (anonymous):

then what do i do next?

OpenStudy (anonymous):

The second Matrix is [3] [2] [1]

OpenStudy (anonymous):

Multiply them First matrices is a 2x3 second is a 3x1 So you answer will be a 2x1

OpenStudy (anonymous):

6+14+6=26 15+8+5=28 Giving you a matrix of [26] [28]

OpenStudy (anonymous):

thanks! can you help me with this one too Use matrices to solve ths system.

OpenStudy (anonymous):

I can kind of help you. \[\left[\begin{matrix}2 & 3 \\ 5 & 8\end{matrix}\right]\left[\begin{matrix}x \\ y \end{matrix}\right]=\left[\begin{matrix}4\\ 11\end{matrix}\right]\]

OpenStudy (anonymous):

\[A^{-1}=\frac{ 1 }{ \det A }\left[\begin{matrix}8 & -3 \\ -5 & 2\end{matrix}\right]\]

OpenStudy (anonymous):

\[A^{1}=\frac{ 1 }{ 16-15 }= \frac{ 1 }{ 1 }=1\left[\begin{matrix}8 & -3 \\ -5 & 2\end{matrix}\right]\]

OpenStudy (anonymous):

\[A^{-1}=\left[\begin{matrix}8 & -3 \\ -5 & 2\end{matrix}\right]\]

OpenStudy (anonymous):

\[\left[\begin{matrix}8 & -3 \\ -5 & 2\end{matrix}\right] \left[\begin{matrix}2 & 3 \\ 5 & 8\end{matrix}\right] \left[\begin{matrix}x \\ y\end{matrix}\right]=\left[\begin{matrix}8 & -3 \\ -5 & 2\end{matrix}\right] \left[\begin{matrix}4 \\11\end{matrix}\right]\]

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 0 \\ 0 & 1\end{matrix}\right] \left[\begin{matrix}x \\ y\end{matrix}\right]= \left[\begin{matrix}-1 \\ 2\end{matrix}\right]\]

OpenStudy (anonymous):

@Best_Mathematician What happens to the matrix \[\left[\begin{matrix}1 & 0 \\ 0 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

@ladydeadpool \[\left[\begin{matrix}x\\ y\end{matrix}\right]=\left[\begin{matrix}-1 \\ 2 \end{matrix}\right]\]

OpenStudy (anonymous):

i really dont know. I never took stats

OpenStudy (anonymous):

can you help me with this last one too? A triangle has vertices (1, 4), (1, 1), and (−3, 1). The triangle is dilated by a scale factor of 2, then translated 5 units up, and then rotated 90° counterclockwise about the origin. What are the vertices of the image of the triangle?

OpenStudy (anonymous):

Yeah hold on a sec. @Best_Mathematician Thanks anyways.

OpenStudy (anonymous):

\[2\left[\begin{matrix}1 & 1 & -3 \\ 4 & 1 & 1\end{matrix}\right]=\left[\begin{matrix}2 & 2 & -6 \\ 8 & 2 & 2\end{matrix}\right]\]Dilation^^

OpenStudy (anonymous):

\[\left[\begin{matrix}2 & 2 & -6 \\ 8 & 2 & 2\end{matrix}\right]+\left[\begin{matrix}0 & 0 & 0 \\ 5 & 5 & 5\end{matrix}\right]=\left[\begin{matrix}2 & 2 & -6\\ 13 & 7 & 7\end{matrix}\right]\] Translation^^

OpenStudy (anonymous):

\[\left[\begin{matrix}0 & -1 \\ -1 & 0\end{matrix}\right]\left[\begin{matrix}2 & 2 & -6 \\ 13 & 7 & 7\end{matrix}\right]=\left[\begin{matrix}-13 & -7 & -7 \\ -2 & -2 & 6\end{matrix}\right]\]Rotation^^

OpenStudy (anonymous):

Final points for A (-13,-2) B (-7,-2) C (-7, 6)

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

No problem!